OpenStudy (anonymous):

which of the following is the equation of a line parallel to the line t=4x+1 passing through the point 5,1 please help :)

OpenStudy (igreen):

I would use a graphing calculator for this. Your answer will be t = 4x - 19 https://www.desmos.com/calculator

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. \(\bbox[5pt, #000000 ,border:5px solid #00FF00]{\bf\huge \color{#00FF00}\sim \! \! \! \! \color{#00FF00}i\it\Huge\color{#00FF00}G\color{#00FF00}r\color{#00FF00}e\color{#00FF00}e\color{#00FF00}n\bf\color{#00FF00}ツ}\)

OpenStudy (igreen):

@nick789789

OpenStudy (anonymous):

igreen the answer you came up with isnt one of my choices i have A 4x-y=19 B-4x-y=19 C 4x+y=-19 D 4x+y=19

OpenStudy (igreen):

Those are just in different forms.

OpenStudy (haseeb96):

the slope of the given line is m1=4 and the point is given that is ( 5,1) here is the condition of paralle so the slopes will be same of both the equations m1=m2 so m1=m2=4 here we use the point slope formula (y2-y1)=m(x2-x1) put the point and the slope in this formula and get the answer (y-1)=4(x-5) y-1=4x-20 4x-y-20+1=0 4x-y-19=0 is the equation of required answer of question

OpenStudy (anonymous):

so what would the answer be in the form i need

OpenStudy (igreen):

The answer is A..my answer(y = 4x - 19) is the same as option A(4x - y = 19).

OpenStudy (anonymous):

okay thank you both for your help

OpenStudy (igreen):

No problem :)